Objective: measure the noise floor on the optical table and the floor so we can decide if the table needs better anchoring before swapping in
the larger optical table
The accelerometrs labeled as MC1 ( just north east of IOO chamber floor ) and MC2 ( north east leg of MC2 table floor ) were moved:
MC1 to the floor at the north west leg of optical table.
MC2 is in the north east corner of the optical table
Atm2 was taken after table leg bolts were tighed at 40 ft/lb
The spectrum looks similar to ETMY (Guralp for below 3 Hz ) except the Z direction